OpenStudy (anonymous):

The sum of K consecutive integers is 41. If the least integer is -40, then k?

OpenStudy (anonymous):

Here's a way of thinking about this. If you're adding consecutive integers and the first is -41, then you're adding:\[-41+(-40)+(-39)+(-38)+...=41\]However, at some point, you're going to get into the positive numbers when you're adding.\[-41+(-40)+(-39)+...+(-3)+(-2)+(-1)+(0)+(1)+(2)+(3)+...=41\]At the point where you get to the positive numbers, the negatives and positives will begin to cancel out. This is because -3 + 3 = 0, -2 + 2 = 0, -1 + 1 = 0. Does this help you with the rest of the problem?

OpenStudy (anonymous):

The cutoff part says \(+...=41\)

OpenStudy (phi):

the least number is -40

OpenStudy (misty1212):

oops

OpenStudy (anonymous):

Ok, if the first numebr is -40, just omit the -41 in my equations.

OpenStudy (dan815):

does that equation work

OpenStudy (anonymous):

hmmmmm so it's 41 then?

OpenStudy (anonymous):

Not quite. 41 would be the last term that you would be adding to get the sum to be 41. However, k is the number of numbers you're adding.

OpenStudy (anonymous):

ohhhhh crap .... it's 81

OpenStudy (anonymous):

-40 to 0 is 41 numbers. 1 to 41 is 41 numbers. So, 41 + 41 should be what k is. Is this making sense to you?

OpenStudy (freckles):

\[\sum_{i=1}^k(-40+i)=41 \\ -40k+\frac{k(k+1)}{2}=41\]

OpenStudy (freckles):

it is a quadratic equation

OpenStudy (anonymous):

hmmm you know what i do is i add 41 to -40 and then multiply by 81

OpenStudy (freckles):

I think I should have started it at i=0

OpenStudy (anonymous):

The key here is that they are "consecutive" integers. That's why just multiplying like that doesn't work.

OpenStudy (freckles):

\[-40(k+1)+\frac{k(k+1)}{2}=41\]

OpenStudy (anonymous):

crap so it's 82

OpenStudy (freckles):

almost

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

@freckles I got 82 as well. Might've had an error though.

OpenStudy (freckles):

you said 81 earlier that is right

OpenStudy (freckles):

that is what you get when you solve that quadratic equation above \[\sum_{i=0}^k (-40+i)\] this does mean -40+(-39)+(-38)+...+(-40+k) or if you used those formula things you can write that sum thingy as \[-40(k+1)+\frac{k(k+1)}{2} \text{ which we want to this output } 41 \\ \text{ so we need \to solve } \\ -40(k+1)+\frac{k(k+1)}{2}=41\]

OpenStudy (freckles):

\[\frac{-80(k+1)}{2}+\frac{k(k+1)}{2}=41 \\ \frac{(k+1)}{2}(-80+k)=41 \\ (k+1)(-80+k)=82 \\ k^2-79k-80=82 \\ k^2-79k-80-82=0 \\ k^2-79k-162=0 \\ (k+2)(k-81)=0\]

OpenStudy (freckles):

only one of these k's make sense

OpenStudy (dan815):

82 is right i believe

OpenStudy (dan815):